OpenStudy (anonymous):

Use elimination to solve the system 4x = 3(1 − y) 3y = 4(8 − x)

OpenStudy (anonymous):

probably want to start with \[4x=3-3y\] \[3y=32-4x\]

OpenStudy (anonymous):

then maybe \[4x+3y=3\] \[4x+3y=32\] and then stop because your lines are parallel

OpenStudy (anonymous):

then times the whole bottom equation by -1? so u can elliminate?

OpenStudy (anonymous):

before you do anything, look at see that the left hand side of both lines is identical

OpenStudy (anonymous):

and since the right hand side is not (you have 3 and 38) this means that the lines are parallel, so they do not intersect. no point in going further

OpenStudy (anonymous):

if you multiply the second one by -1 and then add both the x and the y terms will drop out, leaving you with 0 = -29 which means no solution

OpenStudy (anonymous):

what if there isnt two numbers that are the same? like this one 4x + 5y = 3 16x − 15y = 5

OpenStudy (anonymous):

\[4x+5y=3\] \[16x-15y=5\] they are not parallel

OpenStudy (anonymous):

for this one you can multiply the top one across by 3 and add, and the y will drop out but not the x